# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ Appendix A Example: Prisoner’s Dilemma Dynamics
We remark that Example 7 illustrates a prisoner’s dilemma structure. Table 1 presents the normal form game induced by $f_{0.8}$ and quality profile $(10,10)$. It considers two possible actions for each creator: To (fully) share or not to share. Note that Proposition 2 ensures that any other sharing decision is weakly dominated by those options. Interestingly, the no-sharing strategies $s_{1}=0$ and $s_{2}=0$ are dominated. In the classic prisoner’s dilemma, players are better off if they cooperate, but each has an individual incentive to defect. Example 7 presents the analogous phenomenon, where the players are better off if they don’t share their content with GenAI, but each has an incentive to deviate and share, thereby reducing overall utility.
@@ Appendix B Omitted Proofs from Section 3
This section is divided into two parts. The first part presents and proves all the auxiliary lemmas necessary for the proofs of the main results presented in Section 3. The second part provides the complete proofs of these main results, utilizing the auxiliary lemmas established in the first part.
@@
For the rest of the paper, we denote $x_{i}^{\max}$ as the unique positive root of the equation $\mu x^{\gamma}=c_{i}(x)$ (this is well defined, as proved in Lemma 10).
@@ B.1 Auxiliary Lemmas for Section 3’s Main Results
Consider a profile $(\mathbf{x},\mathbf{1})$ and some arbitrary creator $i\in[n]$. Let $x_{i}^{\max}$ be the unique positive root of the equation $\mu x^{\gamma}=c_{i}(x)$. Then, any deviation $(x_{i}^{\prime},s_{i}^{\prime})$ by creator $i$ where $x_{i}^{\prime}>x_{i}^{\max}$ yields negative utility.
Proof of Lemma 10.
@@
Case 1: Deviation with full sharing ($s_{i}^{\prime}=1$).
With $s_{i}^{\prime}=1$ and $s_{j}=1$ for $j\neq i$, we obtain via Lemma 9:
@@
Case 2: Deviation with no sharing ($s_{i}^{\prime}=0$).
@@
Defining $g(x)=c_{i}(x)-\mu x^{\gamma}$, we require $g(x_{i}^{\prime})\leq 0$. We establish that $g(x)$ has exactly one positive root as follows:
First, $g(0)=0$ since $c_{i}(0)=0$. The derivative $g^{\prime}(x)=c_{i}^{\prime}(x)-\mu\gamma x^{\gamma-1}$ satisfies $g^{\prime}(0)<0$ because:
  • If $\gamma<1$, then $g^{\prime}(0)=c_{i}^{\prime}(0)-\lim_{x\to 0^{+}}\mu\gamma x^{\gamma-1}=0-\infty=-\infty$.
  • If $\gamma=1$, then $g^{\prime}(0)=c_{i}^{\prime}(0)-\mu\gamma=0-\mu\gamma<0$.
This indicates that $g(x)$ initially decreases for small positive $x$.
The second derivative $g^{\prime\prime}(x)=c_{i}^{\prime\prime}(x)-\mu\gamma(\gamma-1)x^{\gamma-2}$ is positive for all $x>0$ because $c_{i}^{\prime\prime}(x)>0$ (strict convexity of the cost function) and $\mu\gamma(\gamma-1)x^{\gamma-2}\leq 0$ (since $\gamma\leq 1$). This implies $g(x)$ is strictly convex, with a strictly increasing first derivative.
Since $g^{\prime}(0)<0$ and $g^{\prime}(x)$ is strictly increasing, $g^{\prime}(x)$ crosses zero exactly once for positive $x$, indicating $g(x)$ has exactly one critical point (a minimum). Combined with $\lim_{x\to\infty}g(x)=\infty$ (as $c_{i}(x)$ grows faster than $\mu x^{\gamma}$ for large $x$), this means $g(x)$ crosses the x-axis at most twice: Once at $x=0$ and once at some positive value.
Therefore, there exists a unique positive root $x_{i}^{\max}$ such that $g(x_{i}^{\max})=0$, and for any $x>x_{i}^{\max}$, we have $g(x)>0$.
Consequently, any deviation $(x_{i}^{\prime},s_{i}^{\prime})$ with $x_{i}^{\prime}>x_{i}^{\max}$ yields negative utility for creator $i$. This completes the proof of Lemma 10. ∎
@@
Step 1: Collecting terms associated with each $x_{r}$.
@@
Step 2: Simplifying the cross-terms.
@@
Step 3: Computing the total coefficient.
@@
Step 4: Proving non-negativity through completing the square.
@@
This completes the proof of Lemma 11. ∎
@@
Proof of Lemma 12.
Consider creator $i$ with arbitrary strategy profiles $(\mathbf{x}_{-i},\mathbf{s}_{-i})$ for the other creators. We demonstrate that zero quality is strictly dominated by some positive quality choice, regardless of the sharing decision.
First, observe that a strategy $(x_{i}=0,s_{i})$ yields zero utility for any sharing level $s_{i}\in[0,1]$. Consider an alternative strategy $(x_{i}=\varepsilon,s_{i}=0)$ for some small $\varepsilon>0$. With this choice:
  • Creator $i$ does not share content: $x_{i}s_{i}=0$ and GenAI quality remains $Q_{\text{AI}}=\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}$.
  • Total quality increases to $\left\|\mathbf{x}\right\|=\varepsilon+\left\|\mathbf{x}_{-i}\right\|$, and the platform traffic becomes $T(\mathbf{x})=\mu(\varepsilon+\left\|\mathbf{x}_{-i}\right\|)^{\gamma}$.
Since $s_{i}=0$, creator $i$ receives no share of AI-driven revenue. Therefore, utility is derived solely from direct traffic:
@@
$U_{i}(\varepsilon,0)=\mu(\varepsilon+\left\|\mathbf{x}_{-i}\right\|)^{\gamma}\frac{\varepsilon}{(\varepsilon+\left\|\mathbf{x}_{-i}\right\|)+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}-c_{i}(\varepsilon).$
Let $Y=\left\|\mathbf{x}_{-i}\right\|$ and $Z=\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}$ for notational simplicity. The utility can be rewritten as:
$U_{i}(\varepsilon,0)=\mu(\varepsilon+Y)^{\gamma}\frac{\varepsilon}{\varepsilon+Y+Z}-c_{i}(\varepsilon).$
To establish that this deviation is profitable, we examine the marginal utility at $\varepsilon=0^{+}$:
@@
$\left.\frac{dU_{i}(\varepsilon,0)}{d\varepsilon}\right|_{\varepsilon=0^{+}}=\left.\frac{d}{d\varepsilon}\left[\mu(\varepsilon+Y)^{\gamma}\frac{\varepsilon}{\varepsilon+Y+Z}-c_{i}(\varepsilon)\right]\right|_{\varepsilon=0^{+}}.$
By assumption, $c_{i}^{\prime}(0)=0$, so it suffices to differentiate the first term. Using the product and quotient rules, we have:
@@
$\frac{d}{d\varepsilon}\left[(\varepsilon+Y)^{\gamma}\frac{\varepsilon}{\varepsilon+Y+Z}\right]$
$=\gamma(\varepsilon+Y)^{\gamma-1}\frac{\varepsilon}{\varepsilon+Y+Z}+(\varepsilon+Y)^{\gamma}\frac{(\varepsilon+Y+Z)-\varepsilon}{(\varepsilon+Y+Z)^{2}}$
$=\gamma(\varepsilon+Y)^{\gamma-1}\frac{\varepsilon}{\varepsilon+Y+Z}+(\varepsilon+Y)^{\gamma}\frac{Y+Z}{(\varepsilon+Y+Z)^{2}}.$
Evaluating at $\varepsilon=0$:
  • If $Y+Z>0$, the first term vanishes and the second becomes $\frac{Y^{\gamma}}{Y+Z}$. Thus,
    $\left.\frac{dU_{i}(\varepsilon,0)}{d\varepsilon}\right|_{\varepsilon=0^{+}}=\mu\frac{Y^{\gamma}}{Y+Z}>0.$
  • If $Y=Z=0$, then $U_{i}(\varepsilon,0)=\mu\varepsilon^{\gamma}-c_{i}(\varepsilon)$, so
    $\left.\frac{dU_{i}(\varepsilon,0)}{d\varepsilon}\right|_{\varepsilon=0^{+}}=\lim_{\varepsilon\to 0^{+}}\mu\gamma\varepsilon^{\gamma-1},$
    which is +$\infty$+ ∞ if $\gamma<1$ and $\mu$ if $\gamma=1$, both strictly positive.
In both cases, the marginal utility at $\varepsilon=0^{+}$ is strictly positive. By continuity, there exists some sufficiently small $\varepsilon>0$ for which $U_{i}(\varepsilon,0)>0=U_{i}(0,s_{i})$. This renders any strategy with $x_{i}=0$ non-rational, regardless of the sharing decision $s_{i}$ or the strategies of other creators. This completes the proof of Lemma 12. ∎
@@ B.2 Proofs of Section 3’s Main Results
$U_{i}(\mathbf{x},(s_{i},\mathbf{s}_{-i});f_{\rho})=T\cdot\frac{x_{i}}{D(s_{i})}+T\cdot\frac{\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})}{D(s_{i})}\cdot\rho\frac{x_{i}s_{i}}{x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}-c_{i}(x_{i}).$
We can further simplify this expression as follows:
$U_{i}(s_{i})$ $=T\cdot\frac{x_{i}}{D(s_{i})}+T\cdot\frac{\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})}{D(s_{i})}\cdot\rho\frac{x_{i}s_{i}}{x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}-c_{i}(x_{i})$
@@
Notice that this simplified expression is consistent with Equation 3 when $s_{i}=0$.
To determine the optimal $s_{i}$, we compute the derivative of $U_{i}(s_{i})$ with respect to $s_{i}$. Since $T$ and $c_{i}(x_{i})$ are constant with respect to $s_{i}$, we focus on differentiating the bracketed term. Its derivative is given by the quotient rule:
$\frac{d}{ds_{i}}\left(\frac{x_{i}(1+\alpha\rho s_{i})}{D(s_{i})}\right)$ $=\frac{(x_{i}\alpha\rho)\cdot D(s_{i})-x_{i}(1+\alpha\rho s_{i})\cdot(\alpha x_{i})}{[D(s_{i})]^{2}}$
$=\frac{\alpha x_{i}}{[D(s_{i})]^{2}}\left[\rho D(s_{i})-x_{i}(1+\alpha\rho s_{i})\right]$
$=\frac{\alpha x_{i}}{[D(s_{i})]^{2}}\left[\rho\left((x_{i}+\left\|\mathbf{x}_{-i}\right\|)+\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})\right)-x_{i}(1+\alpha\rho s_{i})\right]$
$=\frac{\alpha x_{i}}{[D(s_{i})]^{2}}\left[\rho(x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})-x_{i}\right].$
Therefore, the derivative of the utility function is:
$\frac{\partial U_{i}(s_{i})}{\partial s_{i}}=T\cdot\frac{\alpha x_{i}}{[D(s_{i})]^{2}}\left[\rho(x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})-x_{i}\right].$
Since all factors outside the brackets ($T$, $\alpha$, $x_{i}$, and $[D(s_{i})]^{2}$) are strictly positive, the sign of this derivative depends solely on the term in brackets. Specifically, $\frac{\partial U_{i}(s_{i})}{\partial s_{i}}>0$ if and only if:
$\rho(x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})-x_{i}>0$
$\Rightarrow$ $\rho>\frac{x_{i}}{x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}=\tau_{i}.$
@@
This completes the proof of Proposition 2
@@
For enforced sharing ($s_{i}=1$ for all $i$), our model bears structural similarities to Yao et al. [52]. An alternative route to proving this proposition would be to use a reduction to their model, which holds only for enforced sharing. Nevertheless, we provide a direct, self-contained proof tailored to our setting. We stress that any of the other results in this paper, which are strongly connected to the sharing aspect of our model, necessitate fundamentally different arguments than those in Yao et al. [52]. We establish existence and uniqueness by demonstrating that the enforced sharing game is monotone and the utility functions are strictly concave in each player’s action.
Step 1: Utility structure and gradient computation.
@@
Lemma 9 asserts that, under enforced sharing, creator $i$’s utility is:
$U_{i}(\mathbf{x},\mathbf{1};f_{\rho})=\frac{\mu(1+\alpha\rho)}{1+\alpha}\cdot(\|\mathbf{x}\|)^{\gamma-1}\cdot x_{i}-c_{i}(x_{i}).$
@@
Let $K=\frac{\mu(1+\alpha\rho)}{1+\alpha}>0$. Taking the partial derivative with respect to $x_{i}$:
$\frac{\partial U_{i}}{\partial x_{i}}$ $=K\frac{\partial}{\partial x_{i}}\left[\|\mathbf{x}\|^{\gamma-1}x_{i}\right]-c_{i}^{\prime}(x_{i})$
$=K\left[(\gamma-1)\|\mathbf{x}\|^{\gamma-2}\frac{\partial\|\mathbf{x}\|}{\partial x_{i}}x_{i}+\|\mathbf{x}\|^{\gamma-1}\right]-c_{i}^{\prime}(x_{i}).$
Since $\frac{\partial\|\mathbf{x}\|}{\partial x_{i}}=1$ and $\|\mathbf{x}\|=x_{i}+\|\mathbf{x}_{-i}\|$:
$\frac{\partial U_{i}}{\partial x_{i}}$ $=K\left[(\gamma-1)\|\mathbf{x}\|^{\gamma-2}x_{i}+\|\mathbf{x}\|^{\gamma-1}\right]-c_{i}^{\prime}(x_{i})$
$=K\|\mathbf{x}\|^{\gamma-2}\left[(\gamma-1)x_{i}+\|\mathbf{x}\|\right]-c_{i}^{\prime}(x_{i})$
$=K\|\mathbf{x}\|^{\gamma-2}\left[(\gamma-1)x_{i}+x_{i}+\|\mathbf{x}_{-i}\|\right]-c_{i}^{\prime}(x_{i})$
$=K\|\mathbf{x}\|^{\gamma-2}\left[\gamma x_{i}+\|\mathbf{x}_{-i}\|\right]-c_{i}^{\prime}(x_{i}).$
@@
Let $v_{i}(\mathbf{x})=\frac{\partial U_{i}}{\partial x_{i}}$ denote this gradient, and define:
$A_{i}(\mathbf{x})=K\|\mathbf{x}\|^{\gamma-2}\left[\gamma x_{i}+\|\mathbf{x}_{-i}\|\right].$
Then $v_{i}(\mathbf{x})=A_{i}(\mathbf{x})-c_{i}^{\prime}(x_{i})$.
Step 2: Strategy space consideration.

### v2

@@ A.1 ESE Computation
We solve this system numerically using scipy.optimize’s root-finding routines. Under our cost assumptions (in particular, $\lim_{x\to 0^{+}}c_{i}^{\prime}(x)=0$), interior solutions are guaranteed, and the strict concavity ensures that any profile with near-zero gradients closely approximates the true ESE. This approach leverages highly optimized numerical solvers while maintaining theoretical soundness. We use a tolerance of $10^{-8}$ for gradient norms in all experiments.
A.2 FSE Stability Check
Algorithm 3 verifies whether a given profile is an FSE by checking if any creator can profitably deviate to not sharing ($s_{i}=0$). This is necessary and sufficient due to Lemma˜2: any profitable deviation is dominated by one with $s_{i}\in\{0,1\}$, and deviations with $s_{i}=1$ are already ruled out by the ESE property. For each creator, we solve a one-dimensional optimization problem to find their best response when withholding content. The profile is deemed stable only if no creator benefits from such deviation beyond the tolerance threshold $\varepsilon_{tol}$.
Algorithm 3 FSE Stability Check via $s_{i}=0$ Deviation
3:for $i\leftarrow 1$ to $n$ do
4:  Calculate utility at candidate profile: $U_{i}^{\star}\leftarrow U_{i}(x_{i}^{\star},1;x_{-i}^{\star},\mathbf{1}_{-i})$
5:  Calculate rational quality upper bound for deviation: $x_{i}^{max}\leftarrow(\mu/a_{i})^{1/(\theta_{i}-\gamma)}$
6:  $U_{dev}(s_{i}=0)\leftarrow\max_{x_{i}^{\prime}\in[0,x_{i}^{max}]}\{U_{i}(x_{i}^{\prime},0;x_{-i}^{\star},\mathbf{1}_{-i})\}$ $\triangleright$ Requires 1D solver
7:  Check stability condition:
8:  if $U_{dev}(s_{i}=0)-U_{i}^{\star}>\varepsilon_{tol}$ then
9:   return False $\triangleright$ Player $i$ has a profitable deviation
A.3 Simulation Pipeline and Computational Resources
For each parameter combination, we initialized 150 independent game instances with randomly sampled cost coefficients. For each instance, we evaluated 100 equally-spaced values of $\rho\in[0,1]$. The computational workflow proceeds as follows: for each $\rho$ value, we compute the unique ESE by solving the first-order conditions using scipy.optimize, leveraging the strict concavity established in our theoretical analysis. Subsequently, Algorithm˜3 verifies whether each computed ESE also constitutes an FSE by checking for profitable deviations to non-sharing strategies. Results are averaged across instances with 95% confidence intervals computed via bootstrapping.
The complete simulation suite was executed on a standard MacBook Pro with 18 GB of RAM and 8 CPU cores. We used parallel processing across 6 cores to accelerate computation. The full set of experiments required approximately 6 hours of computation time for ESE computation and FSE verification across all parameter combinations. Statistical analysis, bootstrapping for confidence intervals, and visualization required less than 1 minute. The implementation uses standard Python scientific computing libraries (NumPy, SciPy) without GPU acceleration.
@@ Appendix B Omitted Proofs from Section 3
This section is divided into two parts. In the first part, we state and prove all the auxiliary lemmas necessary for the proofs of the main results presented in Section 3. The second part provides the complete proofs of these main results.
@@
For the rest of the paper, we denote $x_{i}^{\max}$ as the unique positive solution to $c_{i}(x)=\mu x^{\gamma}$. This value is well-defined because the function $g(x)=c_{i}(x)-\mu x^{\gamma}$ is strictly convex (as $c_{i}$ is strictly convex and $x^{\gamma}$ is concave) and satisfies $g(0^{+})<0$ and $\lim_{x\to\infty}g(x)=\infty$. The boundary conditions imply that $g(x)$ must eventually become positive, ensuring it crosses zero at least once. Meanwhile, strict convexity ensures that the slope is strictly increasing, meaning $g(x)$ can cross the x-axis exactly once.
@@ B.1 Auxiliary Lemmas for Section 3’s Main Results
Consider a profile $(\mathbf{x},\mathbf{1})$ and some arbitrary creator $i\in[n]$. Then, any deviation $(x_{i}^{\prime},s_{i}^{\prime})$ by creator $i$ where $x_{i}^{\prime}>x_{i}^{\max}$ yields negative utility.
Proof of Lemma˜13.
@@
Case 1: Deviation with full sharing ($s_{i}^{\prime}=1$)
With $s_{i}^{\prime}=1$ and $s_{j}=1$ for $j\neq i$, we obtain via Lemma˜12:
@@
Case 2: Deviation with no sharing ($s_{i}^{\prime}=0$)
@@
Consequently, any deviation $(x_{i}^{\prime},s_{i}^{\prime})$ with $x_{i}^{\prime}>x_{i}^{\max}$ yields negative utility for creator $i$. This completes the proof of Lemma˜13. ∎
@@
Step 1: Collecting terms associated with each $x_{r}$
@@
Step 2: Simplifying the cross-terms
@@
Step 3: Computing the total coefficient
@@
Step 4: Proving non-negativity through completing the square
@@
This completes the proof of Lemma˜14. ∎
@@
Proof of Lemma˜15.
Fix creator $i$ and strategies $(\mathbf{x}_{-i},\mathbf{s}_{-i})$. If $x_{i}=0$, then creator $i$ gets no direct-traffic share and no AI-revenue share, and pays cost $c_{i}(0)=0$, hence $U_{i}(0,s_{i})=0$ for every $s_{i}\in[0,1]$.
@@
Next, let $Y=\left\|\mathbf{x}_{-i}\right\|$ and $Z=\alpha\mathbf{x}_{-i}^{\top}\mathbf{s}_{-i}$. Consider the deviation $(x_{i},s_{i})=(\varepsilon,0)$ with $\varepsilon>0$. Then
$U_{i}(\varepsilon,0)=\mu(\varepsilon+Y)^{\gamma}\cdot\frac{\varepsilon}{\varepsilon+Y+Z}-c_{i}(\varepsilon)=\varepsilon\cdot g(\varepsilon)-c_{i}(\varepsilon),$
where $g(\varepsilon)\coloneqq\frac{\mu(\varepsilon+Y)^{\gamma}}{(\varepsilon+Y+Z)}>0$ for $\varepsilon>0$.
By the mean value theorem, for every $\varepsilon>0$ there exists $\xi\in(0,\varepsilon)$ such that $c_{i}(\varepsilon)-c_{i}(0)=c_{i}^{\prime}(\xi)\varepsilon$. Since $\lim_{x\to 0^{+}}c_{i}^{\prime}(x)=0$, we get $\frac{c_{i}(\varepsilon)}{\varepsilon}\to 0\quad\text{as }\varepsilon\to 0^{+}$, i.e., $c_{i}(\varepsilon)=o(\varepsilon)$.
If $Y>0$, then $g$ is continuous at 0, and, for sufficiently small $\varepsilon$, $\left|g(\varepsilon)-g(0)\right|\leq g(0)/2\implies g(\varepsilon)\geq g(0)/2$. Using $c_{i}(\varepsilon)=o(\varepsilon)$, for all sufficiently small $\varepsilon$ also $c_{i}(\varepsilon)\leq(g(0)/4)\varepsilon$, so
$U_{i}(\varepsilon,0)\;\geq\;\varepsilon\cdot\frac{g(0)}{2}-\frac{g(0)}{4}\varepsilon\;=\;\frac{g(0)}{4}\varepsilon\;>\;0.$
If $Y=0$, then $U_{i}(\varepsilon,0)=\mu\varepsilon^{\gamma}-c_{i}(\varepsilon)$. Since $\gamma\in[0,1]$, we have $\varepsilon=o(\varepsilon^{\gamma})$ when $\gamma<1$ (and when $\gamma=1$ we already know $c_{i}(\varepsilon)=o(\varepsilon)$), so in all cases $c_{i}(\varepsilon)=o(\varepsilon^{\gamma})$. Therefore for all sufficiently small $\varepsilon$, $c_{i}(\varepsilon)\leq(\mu/2)\varepsilon^{\gamma}$, giving $U_{i}(\varepsilon,0)\geq(\mu/2)\varepsilon^{\gamma}>0$.
Thus for every $(\mathbf{x}_{-i},\mathbf{s}_{-i})$ and every $s_{i}$, $(x_{i}=0,s_{i})$ is strictly dominated by $(x_{i}=\varepsilon,s_{i}=0)$ for small enough $\varepsilon>0$. Hence, no rational strategy can play $x_{i}=0$. ∎
Lemma 16.
For any power cost instance where $a_{\min},a_{\max}=O(1)$ and every $\rho\in[0,1]$, it holds that $\left\|\mathcal{X}(\rho)\right\|,\left\|\mathcal{X}_{-i}(\rho)\right\|=\Theta(n^{\frac{\theta-1}{\theta-\gamma}})$.
Proof of Lemma˜16.
Consider any creator $i$. Its FOC in the ESE is:
$\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-2}\left[(\gamma-1)\mathcal{X}_{i}(\rho)+\left\|\mathcal{X}(\rho)\right\|\right]=a_{i}\theta(\mathcal{X}_{i}(\rho))^{\theta-1}.$
Denoting $q_{i}=\frac{\mathcal{X}_{i}(\rho)}{\left\|\mathcal{X}(\rho)\right\|}$, we can rearrange the FOC to get:
$\frac{\mu(1+\alpha\rho)}{1+\alpha}\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\left[1-(1-\gamma)q_{i}\right]=a_{i}\theta(\mathcal{X}_{i}(\rho))^{\theta-1}.$ (4)
Step 1: bound $q_{i}$ uniformly
Let $j^{\star}=\text{arg}\min{j\in[n]}q_{j}$. Since $q_{j^{\star}}\leq\frac{1}{n}\leq\frac{1}{2}$, we have:
$1-(1-\gamma)q_{j^{\star}}\geq 1-\frac{1-\gamma}{2}=\frac{1+\gamma}{2}.$
Now, we divide the FOC of any creator $i$, given by Equation˜4, by the FOC of creator $j^{\star}$:
$\frac{1-(1-\gamma)q_{i}}{1-(1-\gamma)q_{j^{\star}}}=\frac{a_{i}\theta(\mathcal{X}_{i}(\rho))^{\theta-1}}{a_{j^{\star}}\theta(\mathcal{X}_{j^{\star}}(\rho))^{\theta-1}}$
Noticing that the LHS is at most $\frac{1}{\frac{1+\gamma}{2}}=\frac{2}{1+\gamma}$, we have:
@@
$\mathcal{X}_{i}(\rho)\leq\left(\frac{2a_{j^{\star}}}{(1+\gamma)a_{i}}\right)^{\frac{1}{\theta-1}}\mathcal{X}_{j^{\star}}(\rho)\leq\left(\frac{2a_{\max}}{(1+\gamma)a_{\min}}\right)^{\frac{1}{\theta-1}}\mathcal{X}_{j^{\star}}(\rho)\eqqcolon C\mathcal{X}_{j^{\star}}(\rho).$ (5)
Notice that this is true for every creator $j$, not just $j^{\star}$. Substituting Equation˜5 back into the definition of $q_{i}$, we get:
$q_{i}=\frac{\mathcal{X}_{i}(\rho)}{\left\|\mathcal{X}(\rho)\right\|}\leq\frac{\max_{j\in[n]}\mathcal{X}_{j}(\rho)}{\max_{j\in[n]}\mathcal{X}_{j}(\rho)+(n-1)\min_{j\in[n]}\mathcal{X}_{j}(\rho)}\leq\frac{C}{C+n-1}\leq\frac{C}{C+1}.$ (6)
Furthermore, we get that for every creator $i$:
$1\geq 1-(1-\gamma)q_{i}\geq 1-(1-\gamma)\frac{C}{C+1}=\frac{1+\gamma C}{C+1}\coloneqq D>0.$ (7)
Step 2: bound $\left\|\mathcal{X}(\rho)\right\|$
Substituting Equation˜7 back into the FOC of any creator $i$ given by Equation˜4, we have:
$a_{i}\theta(\mathcal{X}_{i}(\rho))^{\theta-1}=\Theta\left(\left\|\mathcal{X}(\rho)\right\|^{\gamma-1}\right)\implies\mathcal{X}_{i}(\rho)=\Theta\left(\left\|\mathcal{X}(\rho)\right\|^{\frac{\gamma-1}{\theta-1}}\right).$ (8)
Summing Equation˜8 over all creators $i\in[n]$, we get:
@@
$\sum_{i=1}^{n}\mathcal{X}_{i}(\rho)=\left\|\mathcal{X}(\rho)\right\|=\Theta\left(n\cdot\left\|\mathcal{X}(\rho)\right\|^{\frac{\gamma-1}{\theta-1}}\right)\implies\left\|\mathcal{X}(\rho)\right\|=\Theta\left(n^{\frac{\theta-1}{\theta-\gamma}}\right).$
Additionally, since $q_{i}\leq\frac{C}{C+1}<1$ for every creator $i$ by Equation˜6, it means that $\left\|\mathcal{X}_{-i}(\rho)\right\|=\Theta\left(\left\|\mathcal{X}(\rho)\right\|\right)=\Theta\left(n^{\frac{\theta-1}{\theta-\gamma}}\right)$ as well. This completes the proof of Lemma˜16. ∎
@@ B.2 Proofs of Section 3’s Main Results
$U_{i}(s_{i})$ $=T\cdot\frac{x_{i}}{D(s_{i})}+T\cdot\frac{\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})}{D(s_{i})}\cdot\rho\frac{x_{i}s_{i}}{x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}-c_{i}(x_{i}).$
@@
Notice that this simplified expression is consistent with Equation˜9 when $s_{i}=0$. Next, we compute the derivative of $U_{i}(s_{i})$ with respect to $s_{i}$:
$\frac{\partial U_{i}(s_{i})}{\partial s_{i}}$ $=T\frac{(x_{i}\alpha\rho)\cdot D(s_{i})-x_{i}(1+\alpha\rho s_{i})\cdot(\alpha x_{i})}{[D(s_{i})]^{2}}$
$=T\frac{\alpha x_{i}}{[D(s_{i})]^{2}}\left[\rho D(s_{i})-x_{i}(1+\alpha\rho s_{i})\right]$
$=T\frac{\alpha x_{i}}{[D(s_{i})]^{2}}\left[\rho\left((x_{i}+\left\|\mathbf{x}_{-i}\right\|)+\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})\right)-x_{i}(1+\alpha\rho s_{i})\right]$
$=T\frac{\alpha x_{i}}{[D(s_{i})]^{2}}\left[\rho(x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})-x_{i}\right].$
Since all factors outside the brackets ($T$, $\alpha$, $x_{i}$, and $D(s_{i})]^{2}$) are strictly positive, the sign of this derivative depends solely on the term in brackets. Specifically, $\frac{\partial U_{i}(s_{i})}{\partial s_{i}}>0$ if and only if:
$\rho(x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})-x_{i}>0\implies\rho>\frac{x_{i}}{x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}=\tau_{i}.$
@@
This completes the proof of Lemma˜2
@@
For enforced sharing ($s_{i}=1$ for all $i$), our model bears structural similarities to Yao et al. [50]. An alternative route to proving this proposition would be to use a reduction to their model, which holds only for enforced sharing. Nevertheless, we provide a direct, self-contained proof tailored to our setting. We establish existence and uniqueness by demonstrating that the enforced sharing game is monotone and the utility functions are strictly concave in each player’s action.
Step 1: Utility structure and gradient computation
@@
Lemma˜12 asserts that, under enforced sharing, creator $i$’s utility is:
@@
Let us denote the first term by $R_{i}(\mathbf{x})$, and $K=\frac{\mu(1+\alpha\rho)}{1+\alpha}>0$. Taking the partial derivative of $R_{i}(\mathbf{x})$ w.r.t. $x_{i}$:
$\frac{\partial R_{i}}{\partial x_{i}}$ $=K\cdot\frac{\partial}{\partial x_{i}}\left[(\left\|\mathbf{x}\right\|)^{\gamma-1}\cdot x_{i}\right]$
$=K\left[(\gamma-1)\left\|\mathbf{x}\right\|^{\gamma-2}x_{i}+\left\|\mathbf{x}\right\|^{\gamma-1}\right]$
$=K\left\|\mathbf{x}\right\|^{\gamma-2}\left[(\gamma-1)x_{i}+\left\|\mathbf{x}\right\|\right]$
$=K\left\|\mathbf{x}\right\|^{\gamma-2}\left[\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|\right]$
@@
We denote this gradient by $A_{i}$, and note that $v_{i}\coloneqq\frac{\partial U_{i}}{\partial x_{i}}=A_{i}-c_{i}^{\prime}(x_{i})$.
Step 2: Strategy space consideration